# Continued radio observations of the persistent radio source associated with FRB20190520B provides insights into its origin

### v1

@@ Abstract
Follow-up studies of persistent emission from Fast Radio Burst (FRB) sources are critical for understanding their elusive emission mechanisms and the nature of their progenitors. This work presents new observations of the persistent radio source (PRS) associated with FRB20190520B. We observe a gradual decay in the PRS brightness, which is punctuated by periods of brightening and dimming at both 1.5 GHz and 3 GHz. Furthermore, our low-frequency ($<1$ GHz) observations—the first for this source—reveal evidence of a spectral break which can be attributed to absorption processes. Interpreted within the framework of the magnetar wind nebula model, our data constrain the age of the magnetar progenitor to $53^{+17}_{-10}$ years, broadly consistent with previous work. Assuming the observed 1.5 GHz variability is driven by scintillation, we derive a conservative lower limit for the source’s radius of $>0.52$ pc, for a screen at 10 kpc. The observations presented here challenge the predictions of the previously published best-fit hypernebula model for this source.
Radio transient sources (2008), Transient sources (1851), Radio sources (1358), Radio bursts (1339)
software: CASA (CASA Team et al., 2022), CASA CAPTURE (Kale & Ishwara-Chandra, 2021), emcee (Foreman-Mackey et al., 2013), ChainConsumer (Hinton, 2016), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007)
@@ 1 Introduction
Fast Radio Bursts (FRBs) are highly energetic $\sim$ millisecond duration radio transients (Lorimer et al., 2007). The observations of FRBs using various radio facilities have revealed that some FRBs repeat, while others appear to be non-repeating (Petroff et al., 2022). Among the FRBs that repeat, only two have shown evidence of a periodicity in their activity windows (Petroff et al., 2022, and refs. therein). FRBs experience propagation effects due to interaction and scattering before reaching our detectors. Some propagation effects include: dispersion measure (DM)—the delay in arrival of low frequency radiation compared to the high frequency radiation due to interaction with ionized interstellar medium; rotation measure (RM)—Faraday rotation of the polarization angle due to magnetic field along the line of sight. The measured DM of the FRB can be expressed as a sum of the DM contribution from the Milky Way, the interstellar medium, the intergalactic medium, and the local DM from the host galaxy of the FRB.
@@
The origin of FRBs is still an open question, with many proposed models (Platts et al., 2019). The discovery of an FRB-like burst from galactic magnetar SGR1935+2154 has bolstered interest in magnetar-based models as the leading candidates for explaining FRBs (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020). To date, a total of 865 FRBs have been identified, of which 71 are repeaters (Xu et al., 2023). 96 FRBs have been localized to their host galaxies, providing rich insights into FRB environments (see e.g. Chatterjee et al., 2017; Tendulkar et al., 2021; Marcote et al., 2020; Niu et al., 2022).
Among the localized repeaters, two FRBs—FRB20121102 (Chatterjee et al., 2017) and FRB20190520B (Niu et al., 2022)—are of particular interest given that a compact persistent radio source (PRS) have been observed associated with both, making them key candidates to test FRB progenitor models. A possible PRS has also been reported for FRB 20201124A by Bruni et al. (2024).
@@
FRB20190520B was discovered by Five-hundred-meter Aperture Spherical radio Telescope (FAST) on May 20, 2019 as part of the CRAFT Survey (Li et al., 2018). It has been localized to a star forming dwarf galaxy at a redshift of $z=0.241$ (Niu et al., 2022; Chen et al., 2025). The average DM of bursts from this source is 1207 cm-3pc. Initially, the DM contribution from the host was thought to be exceptionally large $\sim 900$ cm-3pc. However, Lee et al. (2023) investigated the effect of foreground galaxies and estimated a substantially lower host DM. (also see Bhardwaj et al., 2025b). Observations of bursts from this FRB using the Green Bank Telescope (GBT) and Parkes telescope by (Anna-Thomas et al., 2023) have shown that the RM changes sign twice, pointing to a change in the magnetic field component parallel to the propagation, possibly due to a turbulent magnetized screen surrounding the FRB source. Bhandari et al. (2023) used Very-long-baseline interferometry (VLBI) observations to constrain the size of the PRS $<9$ pc and colocated it to within $80$ pc of the FRB. This PRS exhibits a flat spectrum (slope of $-0.4$ at 2020 and $-0.33$ at 2021 epochs) above 1 GHz accompanied by a $\sim 20$% decline in broadband flux between the 2020 and 2021 epochs (Zhang et al., 2023). The PRS has not been detected at frequencies $<1$ GHz in the past.
@@
Some leading models proposed to explain persistent radio emission from FRB sources are— Magnetar Wind Nebula (MWN) model (Margalit & Metzger, 2018) and hypernebula model (Sridhar & Metzger, 2022). The Magnetar Wind Nebula model attributes persistent radio emission to synchrotron emission from magnetized electron-ion interactions with the nebula, powered by flaring activity of a young magnetar. The Hypernebula model explains the persistent emission as the interaction of winds emitted by an accreting black hole with the surrounding medium. Details of the interpretations of these models are discussed in Section 3.
@@ 2 Observations
We observed the PRS associated with FRB20190520B using the wideband receiver backend of upgraded Giant Metrewave Radio Telescope (uGMRT) in three frequency bands: band 3 (central frequency, $\nu_{c}$ = 400 MHz, bandwidth, BW = 200 MHz), band 4 ($\nu_{c}$ = 750  MHz, BW = 400 MHz) and band 5 ($\nu_{c}$ = 1260 MHz, BW = 400 MHz) between 2023 June 16 and 2023 June 20 (Proposal 44_039, PI: Balasubramanian). The band 3 observations under this project were split over two days. We also obtained three epochs of observations each in band 4 and band 5 between 2024 September 05 and 2024 September 21 (Proposal 46_126, PI: Balasubramanian). In addition to our observations, we included archival uGMRT data from proposal 43_054 (PI: Yi Feng) in our analysis. Raw data were downloaded in the FITS format and converted to the CASA (CASA Team et al., 2022) measurement set format. The data were then calibrated and imaged using the automated continuum imaging pipeline CASA-CAPTURE (Kale & Ishwara-Chandra, 2021). For the band 3 data, the two measurement sets were passed onto the automated pipeline separately and then imaged after combining the calibrated measurement sets. Each pipeline run included eight rounds of self-calibration. The band 5 flux density of the persistent source was calculated using the CASA task imfit within a small circular region (of radius 2$\sim$ 2$\times$ 2 × the size of the synthesized beam at band 5) centred at the PRS coordinates. An additional 5% flux density error was added in quadrature to the error obtained from imfit to account for flux calibration errors (see Table 1). No detections were made in Band 3 or in any epoch of Band 4 observations. The upper limit values listed in Table 1 are the 3×3$\times$3 ×RMS value within a large circular region (of radius 20$\sim$ 20$\times$ 20 × the size of the synthesised beam at the respective band) centred at the position of the PRS.
@@ 2.2 VLA observations and archival data
We conducted a single L-band ($\sim$1.5/,GHz) observation of the PRS using the Jansky Very Large Array (VLA) on 2024 August 08. In addition, VLA S-band ($\sim$3 GHz) data of the PRS20190520B field were obtained from the archival data listed under proposal 23A-010 (PI: Yi Feng, taken between 2023 June 07 and 2023 June 24). We obtained the calibrated measurement set when available or started with the raw data. Raw data were calibrated using the automated VLA calibration pipeline  [^1] https://science.nrao.edu/facilities/vla/data-processing/pipeline/CIPL_654. Following calibration, the data were imaged using the automated imaging pipeline  [^2] https://science.nrao.edu/facilities/vla/data-processing/pipeline/VIPL and refined through self-calibration. The best self-calibrated image was selected, and the CASA task imfit was used in a small region around the PRS to estimate the source flux density. An additional 5% flux uncertainty was added in quadrature to account for flux calibration errors.
A summary of all measurements is listed in Table 1. Figure 1 shows the temporal variation of the PRS flux density at different frequency bands, and Figure 2 displays the variation of the spectrum over the epochs. We observe flaring and dimming episodes of the source in the $\approx 1.5$ GHz data, especially in the late 2020 epoch. The possible reason for this observation is discussed in the Section 3.2. For some observations, the image was heavily affected by radio frequency interference (RFI), resulting in an image that was insufficiently clean to use for our analysis (see Table 1 footnote).
@@
Table 1: Summary of radio continuum observations of the persistent radio source associated with FRB20190520B. For non-detections, 3 $\sigma$ upper limits are reported, where $\sigma$ is the RMS flux density measured in a large region of the residual image. Observation dates (in two formats), frequency bands, flux densities and their associated errors, and proposal identifiers/paper references are listed.
@@
2022 Feb 27 59637.208 1.7 210 34
2022 Nov 11 59894.205 1.3 $<$194 GMRT 43_054 (This work)
2022 Nov 29 59912.125 1.3 268 44 GMRT 43_054 (This work)
2022 Dec 27 59940.131 1.3 156 26 GMRT 43_054 (This work)
2023 Jan 24 59968.036 1.3 281 41 GMRT 43_054 (This work)
2023 Feb 21 59996.167 1.3 $<$122 GMRT 43_054 (This work)
2023 Mar 21 60024.962 1.3 $<$144 GMRT 43_054 (This work)
@@
2023 Jun 15 60110.234 3.0 173 11 VLA 23A-010 (This work)
2023 Jun 16 60111.84 0.7 $<$165 GMRT 44_039 (This work)
@@
2023 Jun 18 60113.178 3.0 166 10 VLA 23A-010 (This work)
2023 Jun 18 60113.219 3.0 165 11 VLA 23A-010 (This work)
2023 Jun 19 60114.196 0.3 $<$195 GMRT 44_039 (This work)
@@
2023 Jun 20 60115.214 3.0 151 11 VLA 23A-010 (This work)
2023 Jun 20 60115.255 3.0 143 10 VLA 23A-010 (This work)
2023 Jun 20 60115.297 3.0 152 10 VLA 23A-010 (This work)
@@
2024 Sep 05 60558.585 0.7 $<$155 GMRT 46_126 (This work)
2024 Sep 12 60565.589 1.3 184 16 GMRT 46_126 (This work)
$\dagger$\$$\dagger$\$footnotemark: 2024 Sep 15 60568.594 0.7 GMRT 46_126 (This work)
2024 Sep 21 60574.329 1.3 190 20 GMRT 46_126 (This work)
$\dagger$\$$\dagger$\$footnotemark: 2024 Sep 21 60574.0 0.7 GMRT 46_126 (This work)
footnotetext: Data affected by severe radio frequency interference. Sufficiently clean image was not obtained.
Figure 1: Light curve of the PRS associated with FRB20190520B. The temporal evolution of the radio emission hints at a steady decay of the PRS flux density, with some flaring and dimming seen superimposed on the decay. Previous data plotted here are from Niu et al. (2022) and Zhang et al. (2023).
Figure 2: Spectrum of the PRS associated with FRB20190520B. The upper limits at frequencies below 1 GHz from this work shows the possibility of a spectral break. Previous data plotted here are from Niu et al. (2022) and Zhang et al. (2023).
@@ 3 Physical Interpretations
3.1 Pulsar Wind Nebula Model
Margalit & Metzger (2018) showed that a single expanding magnetized electron-ion nebula powered by a young magnetar can interact with the surrounding medium to produce synchrotron emission. The magnetar is assumed to inject energy into the nebula at a rate given by
@@ 3.1 Pulsar Wind Nebula Model
where $E_{B_{\star}}$ is the free magnetic energy of the magnetar, $\alpha$ is a powerlaw index, and $t_{0}$ is the onset of the active period of the magnetar. The energy injection drives electrons through the surrounding medium, producing the observed persistent emission. At frequencies $\nu$ above the characteristic synchrotron self-absorption frequency $\nu_{SSA}$, the synchrotron luminosity decays as follows (see Margalit & Metzger, 2018, for detailed calculations)
@@
where $A$ is a scaling constant, $t_{\rm{obs}}$ is the time since the first observation of the PRS, $t_{\rm{age}}$ is the age of the persistent source at the time of the first observation. We perform an MCMC fit of Equation 3 using the emcee package (Foreman-Mackey et al., 2013), incorporating all the detections listed in Table 1. The posterior distributions were analysed using ChainConsumer (Hinton, 2016) resulting in the best-fit parameters listed in Table 2. Figure 3 shows the best fit light curves for different frequencies, following the color-code of the data points. The corresponding best fit spectra are shown in Figure 4. The best-fit age of the persistent source, $t_{\rm{age}}$, is $53^{+17}_{-10}$ years, consistent with the allowed age limits derived in Bhandari et al. (2023) and slightly higher than the 16–22 year estimate from Zhao & Wang (2021). The best-fit energy injection power law index is $\alpha=1.75^{+0.09}_{-0.08}$. Excluding the flaring and dimming events observed in late 2022 does not significantly alter the best-fit parameter values.
Figure 3: Light curve of the FRB20190520B PRS, showing the joint spectro-temporal fit using Equation 3. This fit gives an estimate of the age of the persistent source to be $t_{\rm{age}}=53$ years.
Figure 4: Spectrum of the PRS of FRB20190520B showing the joint spectro-temporal fit using Equation 3. The color of the dashed lines represent the different epochs shown in the same colors. The shaded blue region shows the uGMRT upper limits that suggest a break in the spectrum at frequencies $<1\,GHz$.
@@
$\alpha$ uniform (1, $\infty$) $1.75^{+0.09}_{-0.08}$
$t_{\rm{age}}$ (yr) uniform (4,1900) $53^{+17}_{-10}$
$A$ (uJy) uniform (1, $\infty$) $280^{+20}_{-18}$
In addition to the flux decay observed at frequencies $\nu>1$ GHz, the upper limits from our uGMRT observations at $\nu<1$ GHz (from the 2023 and 2024 epochs) suggest a spectral break between 700 MHz and 1.3 GHz (see blue shaded portion in Figure 4). At the observed timescales, such a spectral break is consistent with synchrotron self-absorption. However, the available upper limits are insufficient to constrain the synchrotron self-absorption frequency. Deeper observations at sub-GHz frequencies will be required to quantitatively determine the location of the spectral break.

### v2

@@ Abstract
Follow-up studies of persistent emission from Fast Radio Burst (FRB) sources are critical for understanding their elusive emission mechanisms and the nature of their progenitors. This work presents new observations of the persistent radio source (PRS) associated with FRB20190520B. We observe a gradual decay in the PRS brightness, which is punctuated by periods of brightening and dimming at both 1.5 GHz and 3 GHz. Furthermore, our low-frequency (<1<1 GHz) observations—the first for this source—reveal evidence of a spectral break which can be attributed to absorption processes. Interpreted within the framework of the magnetar wind nebula model, our data constrain the age of the magnetar progenitor to $52^{+16}_{-10}$ years, broadly consistent with previous work. Assuming the observed 1.5 GHz variability is driven by scintillation, we discuss the constraints on the size of the persistent source. The observations presented here challenge the predictions of the previously published best-fit hypernebula model for this source.
software: CASA (2022PASP..134k4501C), CASA CAPTURE (2021ExA....51...95K), emcee (2013PASP..125..306F), ChainConsumer (Hinton2016), Astropy (astropy:2013; astropy:2018; astropy:2022), NumPy (harris2020array), SciPy (2020SciPy-NMeth), Matplotlib (Hunter:2007)
@@ 1 Introduction
Fast Radio Bursts (FRBs) are highly energetic $\sim$ millisecond duration radio transients (2007Sci...318.. [^7] L). The observations of FRBs using various radio facilities have revealed that some FRBs repeat, while others appear to be non-repeating (2022A&ARv..30....2P). Among the FRBs that repeat, only two have shown evidence of a periodicity in their activity windows (2022A&ARv..30....2P, and refs. therein). FRBs experience propagation effects due to interaction and scattering before reaching our detectors. Some propagation effects include: dispersion measure (DM)—the delay in arrival of low frequency radiation compared to the high frequency radiation due to interaction with ionized medium; rotation measure (RM)—Faraday rotation of the polarization angle due to magnetic field along the line of sight. The measured DM of the FRB can be expressed as a sum of the DM contribution from the Milky Way, the interstellar medium, the intergalactic medium, and the local DM from the host galaxy of the FRB.
@@
The origin of FRBs is still an open question, with many proposed models (2019PhR...821....1P). The discovery of an FRB-like burst from galactic magnetar SGR1935+2154 has bolstered interest in magnetar-based models as the leading candidates for explaining FRBs (2020Natur.587...59B; 2020Natur.587...54C). To date, a total of $\approx 1000$ FRB sources have been identified, of which $10\%$ are repeaters (2021ApJS..257...59C; 2024ApJ...969..145C; 2023Univ....9..330X; universe11070206). Less than a 100 FRBs have been localized to their host galaxies, providing rich insights into FRB environments (see e.g. 2017Natur.541...58C; 2021ApJ...908L..12T; 2020Natur.577..190M; 2022Natur.606..873N). Among the population of repeating FRBs, a handful of sources are of particular interest given that a compact persistent radio source (PRS) have been observed associated with them—FRB20121102 (2017Natur.541...58C), FRB20190520B (2022Natur.606..873N), FRB20201124A (2024Natur.632.1014B; 2025ApJ...993..234B) and FRB20190417A (2025arXiv250905174M) and FRB20181030A (2024ApJ...976..199I).
@@
FRB20190520B was discovered by the Five-hundred-meter Aperture Spherical radio Telescope (FAST) on May 20, 2019 as part of the CRAFT Survey (2018IMMag..19..112L). It has been localized to a star-forming dwarf galaxy at a redshift of $z=0.241$ (2022Natur.606..873N; 2025ApJ...982..203C). The average DM of bursts from this source is 1207 cm-3pc. Initially, the DM contribution from the host was thought to be exceptionally large $\sim 900$ cm-3pc. However, 2023ApJ...954L...7L investigated the effect of foreground galaxies and estimated a substantially lower host DM. (also see Bhardwaj_2025). Observations of bursts from this FRB using the Green Bank Telescope (GBT) and Parkes telescope by (2023Sci...380..599A) have shown that the RM changes sign twice, pointing to a change in the magnetic field component parallel to the propagation, possibly due to a turbulent magnetized screen surrounding the FRB source. 2023ApJ...958L..19B used Very-long-baseline interferometry (VLBI) observations to constrain the size of the PRS <9<9 pc and co-located it to within $80$ pc of the FRB. This PRS exhibits a flat spectrum (slope of 0.4-0.4 at 2020 and 0.33-0.33 at 2021 epochs) above 1 GHz accompanied by a $\sim 20$% decline in broadband flux between the 2020 and 2021 epochs (2023ApJ...959...89Z). The PRS has not been detected at frequencies <1<1 GHz in the past.
@@
Some leading models proposed to explain persistent radio emission from FRB sources are—Magnetar Wind Nebula (MWN) model (2018ApJ...868L...4M) and hypernebula model (2022ApJ...937....5S). The Magnetar Wind Nebula model attributes persistent radio emission to synchrotron emission from magnetized electron-ion interactions with the nebula, powered by flaring activity of a young magnetar. The Hypernebula model explains the persistent emission as the interaction of winds emitted by an accreting black hole with the surrounding medium. Details of the interpretations of these models are discussed in Section 3.
@@ 2 Observations
We observed the PRS associated with FRB20190520B using the wideband receiver backend of upgraded Giant Metrewave Radio Telescope (uGMRT) in three frequency bands: band 3 (central frequency, $\nu_{c}$ = 400 MHz, bandwidth, BW = 200 MHz), band 4 ($\nu_{c}$ = 750  MHz, BW = 400 MHz) and band 5 ($\nu_{c}$ = 1260 MHz, BW = 400 MHz) between 2023 June 16 and 2023 June 20 (Proposal 44_039, PI: Balasubramanian). The band 3 observations under this project were split over two days. We also obtained three epochs of observations each in band 4 and band 5 between 2024 September 05 and 2024 September 21 (Proposal 46_126, PI: Balasubramanian). In addition to our observations, we included archival uGMRT data from proposal 43_054 (PI: Yi Feng) in our analysis. Raw data were downloaded in the FITS format and converted to the CASA (2022PASP..134k4501C) measurement set format. The data were then calibrated and imaged using the automated continuum imaging pipeline CASA-CAPTURE (2021ExA....51...95K). All observations used 3C286 as the flux calibrator. J1543-0757 was used as the phase calibrator for observations under 43_054 and 46_126 while J1558-1409 was used as the phase calibrator for observations under 44_039. For the band 3 data, the two measurement sets (under 44_039) were passed onto the automated pipeline separately and then imaged after combining the calibrated measurement sets. Each pipeline run included eight rounds of self-calibration. The band 5 flux density of the persistent source was calculated using the CASA task imfit within a small circular region (of radius $\sim 2\times$ the size of the synthesized beam at band 5) centered at the PRS coordinates. An additional 5% flux density error was added in quadrature to the error obtained from imfit to account for flux calibration errors (see Table 1). No detections were made in Band 3 or in any epoch of Band 4 observations. The upper limit values listed in Table 1 are the $3\timesRMS$ value within a large circular region (of radius $\sim 20\times$ the size of the synthesized beam at the respective band) centered at the position of the PRS. None of the upper limits are used in the analyses in the section that follows. The observation on 2022 Nov 11 (under GMRT 43_053) shows a possible phase calibration issue, which was identified by imaging the flux calibrator using the same calibration solution (see Figure 5 in appendix for a comparison of the non-detection versus detection of the PRS and the flux calibrator). Additionally, the two observations on 2024 Sep 15 and 2024 Sep 21 (under 46_126) were affected by severe radio frequency interference, not allowing for sufficiently clean images.
@@ 2.2 VLA observations and archival data
We conducted a single L-band ($\sim$1.5 GHz) observation of the PRS using the Jansky Very Large Array (VLA) on 2024 August 08. In addition, VLA S-band ($\sim$3 GHz) data of the PRS20190520B field were obtained from the archival data listed under proposal 23A-010 (PI: Yi Feng, taken between 2023 June 07 and 2023 June 24). We obtained the calibrated measurement set when available or started with the raw data. Raw data were calibrated using the automated VLA calibration pipeline  [^1] https://science.nrao.edu/facilities/vla/data-processing/pipeline/CIPL_654. All observations used 3C286 as the flux calibrator. J1543-0757 was used as the phase calibrator for observations under 23A-010, while J1558-1409 was used as the phase calibrator for observations under 24A-409. Following calibration, the data were imaged using the automated imaging pipeline  [^2] https://science.nrao.edu/facilities/vla/data-processing/pipeline/vipl_666 and refined through self-calibration. The best self-calibrated image was selected, and the CASA task imfit was used in a small region around the PRS to estimate the source flux density. An additional 5% flux uncertainty was added in quadrature to account for flux calibration errors. It may be noted that the data under 23A-010 has already been analyzed using custom software based on the Astronomical Image Processing System [^3] https://www.aips.nrao.edu/index.shtml by 2024ApJ...976..165Y. Here, we repeated the analysis of this data using the CASA VLA calibration pipeline to maintain consistency with the other measurements in this work. There are slight differences in the obtained measurements compared to the ones discussed in 2024ApJ...976..165Y due to the different calibration and imaging tools used here.
A summary of all measurements is listed in Table 1. Figure 1 shows the decay of the overall flux density of the PRS over time. Figure 2 shows the temporal variation of the PRS flux density at different frequency bands, and Figure 3 displays the variation of the spectrum over the epochs. We observe flaring and dimming episodes of the source in the $\approx 1.5$ GHz data, especially in the late 2020 epoch. The possible reason for this observation is discussed in the Section 3.3. For some observations, the image was heavily affected by radio frequency interference (RFI), resulting in an image that was insufficiently clean to use for our analysis (see Table 1 footnote).
@@
Table 1: Summary of radio continuum observations of the persistent radio source associated with FRB20190520B. For non-detections, 3 $\sigma$ upper limits are reported, where $\sigma$ is the RMS flux density measured in a large region of the residual image. Observation dates (in two formats), frequency bands, flux densities and their associated errors, and proposal identifiers/paper references are listed. footnotetext: Probable issue with phase calibration. See non-detection in Figure 5.††††footnotetext: Data affected by severe radio frequency interference. Sufficiently clean image was not obtained.
@@
2022 Nov 29 59912.125 1.3 268 44
2022 Dec 27 59940.131 1.3 156 26
2023 Jan 24 59968.036 1.3 281 41
2023 Feb 21 59996.167 1.3 119 36
2023 Mar 21 60024.962 1.3 280 25
@@
2023 Jun 15 60110.234 3.0 173 11
2023 Jun 16 60111.84 0.7 <<165  $\cdots$ GMRT 44_039 (This work)
@@
2023 Jun 18 60113.178 3.0 166 10
2023 Jun 18 60113.219 3.0 165 11
2023 Jun 19 60114.196 0.3 <<195  $\cdots$ GMRT 44_039 (This work)
@@
2023 Jun 20 60115.214 3.0 151 11
2023 Jun 20 60115.255 3.0 143 10
2023 Jun 20 60115.297 3.0 152 10
@@
2024 Sep 05 60558.585 0.7 <<155  $\cdots$
2024 Sep 12 60565.589 1.3 184 16
[…]
2024 Sep 21 60574.329 1.3 190 20
[…]
@@ 3 Physical Interpretations
3.1 Monotonic trend in the observations
To investigate presence of a monotonic trend in the light curve, we took all the observations listed in Table 1, and scaled the flux densities to 1.5 GHz assuming a spectral index of 0.4-0.4 (as observed in 2023ApJ...959...89Z, for the 2020 epoch). This scaled data was then subjected to the Mann–Kendall test using the pymannkendall package (Hussain2019pyMannKendall). The test yielded a p-value of $1.9\times 10^{-5}$ and a Kendall Tau value of 0.4-0.4, suggesting a significant decreasing trend. Figure 1 shows the light curve with the scaled flux density values along with the linear regression fit (slope =15.3=-15.3) performed with scipy package (2020SciPy-NMeth) showing the decreasing trend. Changing the assumed spectral index to 0.33-0.33 (as observed in 2023ApJ...959...89Z, for the 2021 epoch) results in similar Mann Kendall p-value ($2.8\times 10^{-4}$), Kenadall Tau value (0.4-0.4) and linear regression fit (slope =11.8=-11.8) as above. This establishes a decreasing trend in the total flux density of the PRS as a function of time.
Figure 1: Light curve with all the flux density measurements from Table 1 scaled to 1.5 GHz assuming a spectral index of -0.4. The fit is a linear regression showing a decreasing trend in the overall flux density over time.
3.2 Magnetar Wind Nebula Model
2016MNRAS.461.1498M and 2018ApJ...868L...4M have shown that a single expanding magnetized electron-ion nebula powered by a young magnetar can interact with the surrounding medium to produce synchrotron emission. The magnetar is assumed to inject energy into the nebula at a rate given by
@@ 3.2 Magnetar Wind Nebula Model
where $E_{B_{\star}}$ is the free magnetic energy of the magnetar, $\alpha$ is a powerlaw index, and $t_{0}$ is the onset of the active period of the magnetar. The energy injection drives electrons through the surrounding medium, producing the observed persistent emission. At frequencies $\nu$ above the characteristic synchrotron self-absorption frequency $\nu_{SSA}$, the synchrotron luminosity decays as follows (see 2018ApJ...868L...4M, for detailed calculations)
@@
[…]
Figure 2: Light curve of the FRB20190520B PRS, showing the joint spectro-temporal fit using Equation 3. This fit gives an estimate of the age of the persistent source to be $t_{\rm{age}}=52$ years.
Figure 3: Spectrum of the PRS of FRB20190520B showing the joint spectro-temporal fit using Equation 3. The color of the dashed lines represent the different epochs. The shaded blue region shows the uGMRT upper limits that suggest a break in the spectrum at frequencies <1<1 GHz.
@@
$\alpha$ uniform (1, $\infty$) $1.77^{+0.07}_{-0.08}$
$t_{\rm{age}}$ (yr) uniform (4,1900) $52^{+16}_{-10}$
AA (uJy) uniform (1, $\infty$) $284^{+18}_{-19}$
In addition to the flux decay observed at frequencies $\nu>1$ GHz, the upper limits from our uGMRT observations at $\nu<1$ GHz (from the 2023 and 2024 epochs) suggest a spectral break between 700 MHz and 1.3 GHz (see blue shaded portion in Figure 3). At the observed timescales, such a spectral break is consistent with synchrotron self-absorption. However, the available upper limits are insufficient to constrain the synchrotron self-absorption frequency. Deeper observations at sub-GHz frequencies will be required to quantitatively determine the location of the spectral break. A recent study of the persistent emission from FRB20240114A using GMRT, VLA and MeerKAT data, shows evidence of synchrotron self-absorption peak at $\nu\approx 1.6$ GHz (2025arXiv250114247Z), thereby agreeing with the spectral break below $1.3$ GHz that we observe for the PRS associated with FRB20190520B in this work.